an object of mass 10 kg is projected from ground with speed 40m/s  at an angle 60degree with the horizontal . the rate of change of momentum of object one second after projection in SI unit is:
a.73
b.98
c.170
d.140

Grade:11

1 Answers

Hello
9 years ago
Rate of change of momentum = d(mv)/dt
v = vhorizontal + vvertical 
dv/dt = d(vhorizontal + vvertical) / dt
        =  0 + g   (Since the only force acting is gravity downwards)
Hence d(mv)/dt = m*g = 10*9.8 = 98